# ProRefine: Inference-Time Prompt Refinement with Textual Feedback

### v2

@@ Introduction
We experiment with three models - Llama3.2-1B-instruct, Llama3.2-3B-instruct, and Llama3.1-8B-instruct (Meta 2024) for $LLM_{task}$. The prompts are optimized using Algorithm 1, with Llama3.1-70B-instruct used for feedback generation, prompt optimization, and evaluation. We select the values of hyperparameters $k=10$ and $n=25$ to control the granularity of feedback and duration of optimization. Hyperparameters $k$ and $n$ were fixed based on general preliminary exploration and not tuned per task using benchmark training/validation data.
We compare the performance of our method against the zero-shot Chain-of-Thought (CoT) baseline and TextGrad (Yuksekgonul et al. 2024), and report test accuracy with 95% confidence interval. We choose TextGrad as a baseline because Yuksekgonul et al. (2024) reported performance at par or better than DSPy (Khattab et al. 2024) for prompt optimization on object counting, word sorting, and GSM8k datasets. It is essential to remember that TextGrad is a supervised fine-tuning method that utilizes both the training and validation sets. For TextGrad, we use a comparative setup consisting of a task model to be fine-tuned and Llama3.1-70B-instruct model for feedback generation and backpropagation. The results are shown in Table 1 and Figure 3.
Results
Object Counting
Word Sorting
GSM8K
SVAMP
AQUARAT
Gains over CoT range from 0.920.9-2 percentage points, but declines of 5.4225.4-22 points are also observed. ProRefine exceeds TextGrad on 2 of 3 models, with 2.74.72.7-4.7 percentage point gains, though performance drops of 10.618.110.6-18.1 points are also recorded.
Our results demonstrate that using ProRefine with an optimal verifier significantly improves performance for all tasks, achieving the best results in 13 out of 15 cases, highlighting the critical role of verifier quality. Notably, number of significant improvements increases with larger model sizes.
We also observe that ProRefine enables smaller models, such as Llama3.2-3B-instruct and Llama3.1-8B-instruct, to approach the zero-shot performance of larger models like Llama3.1-8B-instruct and Llama3.1-70B-instruct, respectively. In our experiments using Llama3.1-70B-instruct for $LLM_{task}$, some experiments yielded an accuracy of 1, suggesting potential data contamination. We also experimented using the full chat history at each step, but it failed to produce consistent improvements over our stateless approach. Additionally, we experimented using ProRefine with a comparatively smaller model (Llama3.1-8B-instruct) for $LLM_{feedback}$ and $LLM_{optimizer}$. However, this configuration also failed to consistently yield substantial improvements and, in some cases, even degraded performance. This aligns with prior findings indicating that smaller models are less capable of providing high-quality feedback (Saunders et al. 2022; Bai et al. 2022).
Discussion
@@
Limitations and Future Work

### v3

@@ 5.2 Experimental Setup
We experiment with three models - Llama3.2-1B-instruct, Llama3.2-3B-instruct, and Llama3.1-8B-instruct Meta (2024) for $LLM_{task}$. The prompts are optimized using Algorithm 1, with Llama3.1-70B-instruct used for feedback generation, prompt optimization, and evaluation. We select the values of hyperparameters $k=10$ and $n=25$ to control the granularity of feedback and duration of optimization. Hyperparameters $k$ and $n$ were fixed based on general preliminary exploration and not tuned per task using benchmark training/validation data.
We compare the performance of our method against the zero-shot Chain-of-Thought (CoT) baseline and TextGrad Yuksekgonul et al. (2024), and report test accuracy with 95% confidence interval. We choose TextGrad as a baseline because Yuksekgonul et al. (2024) reported performance at par or better than DSPy Khattab et al. (2024) for prompt optimization on object counting, word sorting, and GSM8k datasets. It is essential to remember that TextGrad is a supervised fine-tuning method that utilizes both the training and validation sets. For TextGrad, we use a comparative setup consisting of a task model to be fine-tuned and Llama3.1-70B-instruct model for feedback generation and backpropagation. The results are shown in Table 1 and Figure 2.
Figure 4: ProRefine example. Given an input query (which in this case has a correct answer: 13) and an initial prompt, the task model ($LLM_{task}$) gives an incorrect answer. ProRefine uses two additional models, $LLM_{feedback}$ and $LLM_{optimizer}$, to iteratively improve the prompt as the $LLM_{task}$ generates its response. Refining the prompt during generation allows the feedback model to target local regions of the response, providing finer-grained feedback than waiting for the response to complete. We also provide an additional example illustrating our approach in Figure 5.
6 Discussion
This work investigates the following research questions.
RQ1
How effectively can textual feedback enhance the performance of LLMs during inference?
RQ2
RQ3
@@ 6 Discussion
7 Limitations and Future Work